OpenStudy (anonymous):

Solve for x 37=5(1.02)^x

OpenStudy (anonymous):

Okay, do you understand how to use logarithms?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, firstly you want to simplify, so divide 5 from both sides

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{37}{5}=1.02^{x}\]

OpenStudy (anonymous):

yup did that and I got 7.4=1.02^x

OpenStudy (anonymous):

Now take the natural logarithm of both sides, so that you have \[\ln(\frac{37}{5})=x * \ln(1.02)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so then x=1.982

OpenStudy (anonymous):

Is that correct

OpenStudy (anonymous):

no

OpenStudy (anonymous):

You want to do \[x=\frac{\ln(\frac{37}{5})}{\ln(1.02)}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Did you do it?

OpenStudy (anonymous):

yes I got 101.071 is that correct

OpenStudy (anonymous):

try it out

OpenStudy (anonymous):

yup thats it thank you

OpenStudy (anonymous):

No problem. :)